OpenStudy (anonymous):

Gary has walked 305 meters, and continues at a rate of 1 meter per second. Mary has walked 113 meters, and continues at a rate of 3 meters per second. In how many seconds will they have traveled the same distance?

OpenStudy (anonymous):

distance = rate x time distance for gary = 305 +1t distance for Mary is 113 + 3t if u want them to have the same distance set the equation to each other and and get t (it is in seconds) 305 +t=113 +3t